Lin Corporation has a single product whose selling price is $140 per unit and whose variable expense is $70 per unit. The company’s monthly fixed expense is $31,900.

Required:

1. Calculate the unit sales needed to attain a target profit of $6,250. (Do not round intermediate calculations.)

2. Calculate the dollar sales needed to attain a target profit of $9,400.
Answer #1

Contribution margin per unit = 140 - 70 = 70

Contribution margin ratio = 70/140 = 50%

Unit sales needed = (Fixed costs +Target profit) /Contribution margin per unit

= (31,900+6,250)/70

= 545 units

Dollar sales needed = (Fixed cost+target profit)/Contribution margin ratio

= (31,900+9,400)/50%

= 82,600